OpenStudy (anonymous):

please help? What is the x-intercept of the graph: 2x + 4y = -10

OpenStudy (klimenkov):

Put y=0 and find x.

OpenStudy (anonymous):

-5,0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this true @klimenkov

OpenStudy (anonymous):

For Ax+By=C The x-intercept is C/A.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

A. (0, -5) B. (-5, 0) c. 0, -5/2 d. -5/-5/2 which one?

OpenStudy (anonymous):

'Constant-over-coefficient' In your equation, 2x + 4y = -10 A=2 B=4 C=-10 C/A = -10/2.

OpenStudy (klimenkov):

B is true.

OpenStudy (anonymous):

the multiple choices are on top